# On an unconditional $\rm GL_{3}$ analog of Selberg’s result

### v1

@@ 2.1. Hecke relations
Firstly, let’s recall the Hecke relations (see Goldfeld [7, pg. 168]):
$\begin{split}A_{F}(m_{1}m_{2},n_{1}n_{2})&=A_{F}(m_{1},n_{1})A_{F}(m_{2},n_{2}),\quad\text{if}\,\,(m_{1}n_{1},m_{2}n_{2})=1,\\ A_{F}(n,1)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\left(\frac{m_{1}d_{0}}{d_{1}},\frac{m_{2}d_{1}}{d_{2}}\right),\\ A_{F}(1,n)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\left(\frac{m_{1}d_{2}}{d_{1}},\frac{m_{2}d_{0}}{d_{2}}\right),\\ A_{F}(n_{1},1)A_{F}(1,n_{2})&=\sum_{d|(n_{1},n_{2})}A_{F}\left(\frac{n_{1}}{d},\frac{n_{2}}{d}\right).\end{split}$
Then we arrive at the following useful result, which was presented by Liu–Liu in [13, Proposition 5.1].
@@ 2.2. A summation formula
We shall need the following summation formula, which is an application of the $\rm GL_{3}$ Kuznetsov trace formula in the version of Buttcane [4, Theorems 2–4].
@@
Let $P=m_{1}n_{1}m_{2}n_{2}\neq 0$. We have for any $\varepsilon>0$ that
$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\overline{A_{F}(m_{1},m_{2})}A_{F}(n_{1},n_{2})\\ &\quad=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\mathcal{H}+O\left((TP)^{\varepsilon}\big{(}TP^{1/2}+T^{3}+TM^{2}P^{\vartheta}\big{)}\right),\end{split}$
where $\vartheta=7/64$ is the best result towards the Ramanujan–Petersson conjecture for Fourier coefficients of $\rm GL_{2}$ Maass cusp forms due to Kim–Sarnak [10].
@@
See Sun–Wang [18, Theorem 7.1]. ∎
We record here the frequently used estimate
@@
for some constant $C>0$ (see Buttcane–Zhou [5, pg. 15]). Taking $m_{1}=n_{1}=m_{2}=n_{2}=1$ in Theorem 2.2 immediately gives
@@
With the assumptions in Theorem 2.2. We have
@@ 2.3. The zero-density estimate
We will need the following weighted version of zero-density estimate in the spectral aspect for the $L$-functions associated to $\rm GL_{3}$ Hecke–Maass cusp forms which was established by Sun–Wang [18, Theorem 1.1].
@@
$\begin{split}N(\sigma,H;F):=\#\big{\{}\rho=1/2+\beta+i\gamma:L(\rho,F)=0,\beta>\sigma,|\gamma|<H\big{\}}\end{split}$
@@
Let $22/\log T<\sigma<1/$ 2. For some sufficiently small $\delta_{1},\,\theta>0$, we have
@@ 3. An approximation of $S_{F}(t)$
In this section we will prove several lemmas and derive an approximation of $S_{F}(t)$. We denote by $\rho=\beta+i\gamma$ a typical zero of $L(s,F)$ inside the critical strip, i.e., $10<\beta<$ 1. For $\operatorname{Re}(s)>1$, we have the Dirichlet series expansion for the logarithmic derivative of $L(s,F)$,
$-\frac{L^{\prime}}{L}(s,F)=\sum_{n=1}^{\infty}\frac{\Lambda(n)C_{F}(n)}{n^{s}},$ (3.1)
@@
Let $x>1$. For $s\neq\rho$, and $s\neq-2m+\mu_{F,j}$, ($m\geq 0$, $j=1,2,3$), we have the following identity
@@
$\Lambda_{x}(n)=\begin{cases}\Lambda(n),\,&\textit{if }\,\,n\leq x,\\ \Lambda(n)\frac{\log^{2}(x^{3}/n)-2\log^{2}(x^{2}/n)}{2\log^{2}x},\,&\textit{if }\,\,x\leq n\leq x^{2},\\ \Lambda(n)\frac{\log^{2}(x^{3}/n)}{2\log^{2}x},\,&\textit{if }\,\,x^{2}\leq n\leq x^{3},\\ 0,\,&\textit{if }\,\,n\geq x^{3}.\\ \end{cases}$
@@
$\frac{1}{2\pi i}\int_{(\alpha)}\frac{y^{s}}{s^{3}}\mathrm{d}s=\begin{cases}\frac{\log^{2}y}{2},\,&\textit{if }\,\,y\geq 1,\\ 0,\,&\textit{if }\,\,0<y\leq 1\end{cases}$
@@
$-\log^{2}x\sum_{n=1}^{\infty}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}=\frac{1}{2\pi i}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v,$
@@
$\begin{split}&\frac{1}{2\pi i}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v\\ =&\frac{L^{\prime}}{L}(s,F)\log^{2}x+\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}+\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}}.\end{split}$
Thus we complete the proof of lemma. ∎
@@
For $s=\sigma+it$, $s^{\prime}=\sigma^{\prime}+it^{\prime}$ such that $1/2\leq\sigma,\sigma^{\prime}\leq 2$, $s\neq\rho$ and $s^{\prime}\neq\rho$, we have
$\operatorname{Im}\left(\frac{L^{\prime}}{L}(s,F)-\frac{L^{\prime}}{L}(s^{\prime},F)\right)=\operatorname{Im}\sum_{\rho}\left(\frac{1}{s-\rho}-\frac{1}{s^{\prime}-\rho}\right)+O(1),$
@@
By the Hadamard factorization theorem of the entire function $\Lambda(s,F)$, we have
$\frac{L^{\prime}}{L}(s,F)=b_{F}+\sum_{\rho}\left(\frac{1}{s-\rho}+\frac{1}{\rho}\right)-\frac{1}{2}\sum_{j=1}^{3}\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{s-\mu_{F,j}}{2}\right)+\frac{3}{2}\log\pi,$
for some $b_{F}\in\mathbb{C}$ with $\operatorname{Re}(b_{F})=-\operatorname{Re}\sum_{\rho}\frac{1}{\rho}$ (see [9, Proposition 5.7 (3)]). Note that for $z\not\in-\mathbb{N}$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=-\gamma+\sum_{k\geq 1}\left(\frac{1}{k}-\frac{1}{k-1+z}\right),$ (3.4)
where $\gamma$ is the Euler constant. Using (3.4), we get that for $\frac{1}{4}\leq u$,
$\operatorname{Im}\frac{\Gamma^{\prime}}{\Gamma}(u+iv)=\sum_{k\geq 1}\frac{v}{(k+u-1)^{2}+v^{2}}\ll 1.$
Combined with the asymptotic formula of the logarithmic derivative of $\Gamma(z)$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=\log z-\frac{1}{2z}+O_{\varepsilon}\left(\frac{1}{|z|^{2}}\right),\quad\text{for}\,\,|\arg z|\leq\pi-\varepsilon,$
we complete the proof of lemma. ∎
@@
$\sigma_{x}=\sigma_{x,F}=\sigma_{x,F,t}:=\frac{1}{2}+2\max_{\rho}\left\{\left|\beta-\frac{1}{2}\right|,\frac{5}{\log x}\right\},$ (3.5)
where $\rho=\beta+i\gamma$ runs through the zeros of $L(s,F)$ for which
@@
$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+O\left(x^{1/4-\sigma/2}\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)\\ &+O\left(x^{1/4-\sigma/2}\log(|t|+M_{F})\right),\end{split}$
@@
$\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}=O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\big{(}\log(|t|+M_{F})\big{)}.$
@@
$\operatorname{Re}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)=\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})).$ (3.6)
Moreover, if $\rho=\beta+i\gamma$ is zero of $L(s,F)$, then $(1-\beta)+i\gamma$ is also a zero of $L(s,F)$. Thus we have
$\begin{split}&\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ =&\frac{1}{2}\left(\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+\sum_{\rho}\frac{\sigma_{x}-(1-\beta)}{(\sigma_{x}-(1-\beta))^{2}+(t-\gamma)^{2}}\right)\\ =&\left(\sigma_{x}-\frac{1}{2}\right)\sum_{\rho}\frac{(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2})}.\end{split}$ (3.7)
@@
If $|\beta-1/2|>\frac{\sigma_{x}-1/2}{2}$, then by (3.5) and (3) we have
@@
$\begin{split}&(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}\\ &=\big{(}(\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big{)}+(t-\gamma)^{2}-2(\beta-1/2)^{2}\\ &\geq\frac{1}{2}\big{(}(\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big{)}+\frac{7}{9}(t-\gamma)^{2}\\ &\geq\frac{1}{4}\big{(}(\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big{)}.\end{split}$ (3.9)
Combining (3.8), (3.9) and (3.7), we have
@@
$\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}\leq\frac{4}{\sigma_{x}-1/2}\left|\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\right|+O\left(\frac{\log(|t|+M_{F})}{\sigma_{x}-1/2}\right).$ (3.10)
On the other hand, by Lemma 3.1 and the bound (1.6), we have
$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+\frac{w_{F}(x,\sigma,t)}{\log^{2}x}\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}^{3/2}}+O\left(\frac{x^{5/14-\sigma}}{\log^{2}x}\right),\end{split}$ (3.11)
@@
[…]
Inserting this bound into (3.11), we have
$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)\leq&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+\frac{4}{5}w^{\prime}_{F}(x,\sigma,t)x^{1/4-\sigma/2}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ &+O\left(x^{1/4-\sigma/2}\log(|t|+M_{F})\right),\end{split}$ (3.13)
@@
$\begin{split}&\left(1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right)\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ =&\,O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\right).\end{split}$
Since $\left|1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right|\geq 1-\left|\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right|\geq\frac{1}{5}$, we have
$\frac{L^{\prime}}{L}(\sigma_{x}+it,F)=O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\right).$ (3.14)
Inserting (3.14) into (3.13) and (3.10), respectively, we complete the proof of lemma. ∎
@@
For $t\neq 0$, and $x\geq 4$, we have
$\begin{split}S_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}\log n}+O\left((\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)},\end{split}$
@@
$\begin{split}\pi S_{F}(t)=&-\int_{1/2}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma\\ =&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma-(\sigma_{x}-1/2)\operatorname{Im}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ &+\int_{1/2}^{\sigma_{x}}\operatorname{Im}\left(\frac{L^{\prime}}{L}(\sigma_{x}+it,F)-\frac{L^{\prime}}{L}(\sigma+it,F)\right)\mathrm{d}\sigma\\ :=&\,J_{1}+J_{2}+J_{3}.\end{split}$

### v2

@@ 2.1. Hecke relations
Firstly, let us recall the following Hecke relations (see Goldfeld [10, pg. 168])
$\begin{split}A_{F}(m_{1}m_{2},n_{1}n_{2})&=A_{F}(m_{1},n_{1})A_{F}(m_{2},n_{2}),\quad\text{if}\,\,(m_{1}n_{1},m_{2}n_{2})=1,\\ A_{F}(n,1)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\bigg(\frac{m_{1}d_{0}}{d_{1}},\frac{m_{2}d_{1}}{d_{2}}\bigg),\\ A_{F}(1,n)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\bigg(\frac{m_{1}d_{2}}{d_{1}},\frac{m_{2}d_{0}}{d_{2}}\bigg),\\ A_{F}(n_{1},1)A_{F}(1,n_{2})&=\sum_{d|(n_{1},n_{2})}A_{F}\bigg(\frac{n_{1}}{d},\frac{n_{2}}{d}\bigg).\end{split}$
Then we arrive at the following useful result, which was presented by Liu and Liu in [15, Proposition 5.1].
@@ 2.2. A summation formula
We need the following summation formula, which is an application of the $\mathrm{GL}_{3}$ Kuznetsov trace formula in the version of Buttcane [6, Theorems 2–4].
@@
Let $P=m_{1}n_{1}m_{2}n_{2}\neq 0$. For any $\varepsilon>0$, we have
$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\overline{A_{F}(m_{1},m_{2})}A_{F}(n_{1},n_{2})\\ &\quad=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\mathcal{H}+O\bigg((TP)^{\varepsilon}\big(TP^{1/2}+T^{3}+TM^{2}P^{\vartheta}\big)\bigg),\end{split}$
where $\vartheta=7/64$, due to Kim and Sarnak [13], is the best record towards the Ramanujan–Petersson conjecture for Fourier coefficients of $\mathrm{GL}_{2}$ Maass cusp forms.
@@
See Sun and Wang [20, Theorem 7.1]. ∎
We quote here the frequently used estimate
@@
for some constant $C>0$ (see Buttcane and Zhou [7, pg. 15]). Taking $m_{1}=n_{1}=m_{2}=n_{2}=1$ in Theorem 2.2 immediately gives
@@
Let the assumptions be as in Theorem 2.2. We have
@@ 2.3. The zero-density estimate
In the context of our proof, the key point lies in the following weighted version of zero-density estimate in the spectral aspect for the $L$-functions associated to $\mathrm{GL}_{3}$ Hecke–Maass cusp forms, which was rigorously established by Sun and Wang in [20, Theorem 1.1].
@@
$\begin{split}N(\sigma,H;F):=\#\big\{\rho=1/2+\beta+\mathrm{i}\gamma:L(\rho,F)=0,\beta>\sigma,|\gamma|<H\big\}\end{split}$
@@
Let $2/\log T<\sigma<1/2$. For some sufficiently small $\delta_{1},\,\theta>0$, we have
@@ 3. An approximation of $S_{F}(t)$
In this section we prove several lemmas and derive an approximation of $S_{F}(t)$. We denote by $\rho=\beta+\mathrm{i}\gamma$ a typical (non-trivial) zero of $L(s,F)$ inside the critical strip, i.e., $0<\beta<1$. For $\operatorname{Re}(s)>1$, we have the Dirichlet series expansion for the logarithmic derivative of $L(s,F)$,
$\frac{L^{\prime}}{L}(s,F)=\sum_{n=1}^{\infty}\frac{\Lambda(n)C_{F}(n)}{n^{s}},$ (3.1)
@@
Let $x>1$. For $s\neq\rho$ and $s\neq-2m+\mu_{F,j}$ ($m\geq 0$, $j=1,2,3$), we have the following identity
@@
[…]
@@
$\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{y^{s}}{s^{3}}\mathrm{d}s=\begin{cases}\frac{\log^{2}y}{2},\,&\textit{if }\,\,y\geq 1,\\ 0,\,&\textit{if }\,\,0<y\leq 1\end{cases}$
@@
$\log^{2}x\sum_{n=1}^{\infty}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}=\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v,$
@@
$\begin{split}&\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v\\ =&\frac{L^{\prime}}{L}(s,F)\log^{2}x+\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}+\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}}.\end{split}$
This completes the proof of the lemma. ∎
@@
For $s=\sigma+\mathrm{i}t$ and $s^{\prime}=\sigma^{\prime}+\mathrm{i}t^{\prime}$ such that $1/2\leq\sigma,\sigma^{\prime}\leq 10$, $s\neq\rho$ and $s^{\prime}\neq\rho$, we have
$\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(s,F)-\frac{L^{\prime}}{L}(s^{\prime},F)\bigg)=\operatorname{Im}\sum_{\rho}\bigg(\frac{1}{s-\rho}-\frac{1}{s^{\prime}-\rho}\bigg)+O(1),$
@@
By the Hadamard factorization theorem for the entire function $\Lambda(s,F)$, we have
$\frac{L^{\prime}}{L}(s,F)=b_{F}+\sum_{\rho}\bigg(\frac{1}{s-\rho}+\frac{1}{\rho}\bigg)-\frac{1}{2}\sum_{j=1}^{3}\frac{\Gamma^{\prime}}{\Gamma}\bigg(\frac{s-\mu_{F,j}}{2}\bigg)+\frac{3}{2}\log\pi,$
for some $b_{F}\in\mathbb{C}$ with $\operatorname{Re}(b_{F})=-\operatorname{Re}\sum_{\rho}\frac{1}{\rho}$ (see [12, Proposition 5.7 (3)]). Note that for $z\not\in$-$\mathbb{N}$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=-\gamma+\sum_{k\geq 1}\bigg(\frac{1}{k}-\frac{1}{k-1+z}\bigg),$ (3.4)
where $\gamma$ is the Euler constant. Using (3.4), we get that for $\frac{1}{14}\leq u\leq\frac{145}{28}$ and $v\ll 1$,
$\operatorname{Im}\frac{\Gamma^{\prime}}{\Gamma}(u+\mathrm{i}v)=\sum_{k\geq 1}\frac{v}{(k+u-1)^{2}+v^{2}}\ll 1.$
Combining with the asymptotic formula of the logarithmic derivative of $\Gamma(z)$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=\log z-\frac{1}{2z}+O_{\varepsilon}\bigg(\frac{1}{|z|^{2}}\bigg),\quad\text{for}\,\,|\arg z|\leq\pi-\varepsilon,$
we complete the proof of the lemma. ∎
@@
$\sigma_{x}=\sigma_{x,F}=\sigma_{x,F,t}:=\frac{1}{2}+2\max_{\rho}\bigg\{\bigg|\beta-\frac{1}{2}\bigg|,\frac{5}{\log x}\bigg\},$ (3.5)
where $\rho=\beta+\mathrm{i}\gamma$ runs through the zeros of $L(s,F)$ for which
@@
$\begin{split}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+O\bigg(x^{1/4-\sigma/2}\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)\\ &+O\bigg(x^{1/4-\sigma/2}\log(|t|+M_{F})\bigg),\end{split}$
@@
$\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}=O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\big(\log(|t|+M_{F})\big).$
@@
$\operatorname{Re}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)=\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})).$ (3.6)
Moreover, if $\rho=\beta+\mathrm{i}\gamma$ is zero of $L(s,F)$, then $(1-\beta)+\mathrm{i}\gamma$ is also a zero of $L(s,F)$ by the functional equation (1.7). Thus we have
$\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}$
$=$ $\,\frac{1}{2}\bigg(\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+\sum_{\rho}\frac{\sigma_{x}-(1-\beta)}{(\sigma_{x}-(1-\beta))^{2}+(t-\gamma)^{2}}\bigg)$
$=$ $\,\bigg(\sigma_{x}-\frac{1}{2}\bigg)\sum_{\rho}\frac{(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2})}.$ (3.7)
@@
If $|\beta-1/2|>\frac{\sigma_{x}-1/2}{2}$, then by (3.5) we have
@@
$(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}$
$=\big((\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big)+(t-\gamma)^{2}-2(\beta-1/2)^{2}$
$\geq\frac{1}{2}\big((\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big)+\frac{7}{9}(t-\gamma)^{2}$
$\geq\frac{1}{4}\big((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big).$ (3.9)
Combining (3.8), (3) and (3), we get
@@
$\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}\leq\frac{4}{\sigma_{x}-1/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg(\frac{\log(|t|+M_{F})}{\sigma_{x}-1/2}\bigg).$ (3.10)
On the other hand, by Lemma 3.1 and the estimate (1.6), we have, for $\sigma\geq\sigma_{x}$,
$\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=$ $-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+\frac{w_{F}(x,\sigma,t)}{\log^{2}x}\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}$
$+O\bigg(\frac{x^{5/14-\sigma}}{\log^{2}x}\bigg),$ (3.11)
@@
$\begin{split}&\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}\\ \leq&\,2\log x\sum_{\rho}\frac{x^{1/4-\sigma/2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}\\ \leq&\,8\log x\frac{x^{1/4-\sigma/2}}{\sigma_{x}-1/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg(\frac{(\log x)x^{1/4-\sigma/2}\log(|t|+M_{F})}{\sigma_{x}-1/2}\bigg)\\ \leq&\,\frac{4}{5}(\log x)^{2}x^{1/4-\sigma/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg((\log x)^{2}x^{1/4-\sigma/2}\log\big(|t|+M_{F}\big)\bigg).\end{split}$
Inserting this bound into (3), we have
$\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=$ $-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+\frac{4}{5}w^{\prime}_{F}(x,\sigma,t)x^{1/4-\sigma/2}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)$
$+O\bigg(x^{1/4-\sigma/2}\log(|t|+M_{F})\bigg),$ (3.13)
@@
$\begin{split}&\bigg(1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\bigg)\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\\ =&\,O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\bigg).\end{split}$
Since $\Big|1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\Big|\geq 1-\Big|\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\Big|\geq\frac{1}{5}$, we have
$\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)=O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\bigg).$ (3.14)
Then the claimed results follow from (3) and (3.14), and from (3.10) and (3.14), respectively. ∎
@@
For $t\neq 0$ and $x\geq 4$, we have
$\begin{split}S_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}\log n}+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big),\end{split}$
@@
$\begin{split}\pi S_{F}(t)=&-\int_{1/2}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma\\ =&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma-(\sigma_{x}-1/2)\operatorname{Im}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\\ &+\int_{1/2}^{\sigma_{x}}\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)-\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\bigg)\mathrm{d}\sigma\\ :=&\,J_{1}+J_{2}+J_{3}.\end{split}$